Use the References to access important values if needed for this question. The vapor pressure of liquid octane, CgH1s, is 100. mm Hg at 339 K in a closed, evacuated 544 ml container at a temperature of 339 K. It is found that all of the CgHis is in the vapor phase and that the pressure is 61.0 mm Hg. If the volume of the container is reduced to 345 ml at the following statements are correct? constant temperature, which of Choose all that apply No condensation will occur. O The pressure in the container will be 100. mm Hg. O Some of the vapor initially present will condense. Only octane vapor will be present. D Liquid octane will be present 7 more group attempts remaining Retry Entire Group Submit Answer
Review Topics] Use the References to access important values if needed for this question. The vapor pressure of liquid propyl amine, C3H,NEH2, is 40.0 mm Hg at 257 K. A sample of C3H-NH2 is placed in a closed, evacuated container of constant volume at a temperature of 377 K. It is found that all of the C3H-NH, is in the vapor phase and that the pressure is 49.0 mm Hg. If the temperature in the container is reduced to 257 K, which of the following statements are correct? Choose all that apply. Only propyl amine vapor will be present. The pressure in the container will be 40.0 mm Hg Liquid propyl amine will be present. Some of the vapor initially present will condense. No condensation will occur 7 more group attempts remaining Retry Entire Group Subrmit Answer
Answer #1

1) The pressure at which the system arrives when the volume decreases is calculated:

P2 = P1 * V1 / V2 = 61 mmHg * 544 mL / 345 mL = 96.19 mmHg

The pressure is still less than the vapor pressure at that temperature.

No condensation occurs, only octane vapor is present.

2) The pressure at which the system arrives when the temperature is lowered is calculated:

P2 = P1 * T2 / T1 = 49 mmHg * 257 K / 377 K = 33.40 mmHg

The pressure is less than the vapor pressure at that temperature.

Liquid amine is present, The vapor contained initially condensed.

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.